AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 218     Entry time: Mon Feb 21 12:10:47 2022
Author: RSS, JG 
Subject: Alpha background run50 

Beam OFF. Run started 20/2/22 18:42

System checck OK, except white rabbit all fail

Temperatures OK

Leakage current 3.59 uA.

potentiometer = 20.28 mm

 

11:36 OH noticed a channel in aida03 was running hot. Did an asic load check and rate returned to normal.

          Online monitor also restarted.

 

15:06 Investigating the high rate in aida03 ASIC01 channel 11.
          Looking at the waveforms it seems channel 11 is consistently different.
          Attachment 1 shows a typical waveform for it

 

Attachment 1: 220221_1506_aida03_1*W.png  237 kB  Uploaded Mon Feb 21 14:09:02 2022  | Hide | Hide all
Cannot create thumbnail, please check ImageMagick installation
ELOG V3.1.4-unknown